Calculate the price elasticity of demand, if there is a decline in quantity from 55 to 45 a


week and an increase in price from $12 to $20.


1
Expert's answer
2020-08-05T16:32:09-0400

Using the arc elasticity formula below:

"\text{Price elasticity of demend}=\dfrac{\dfrac{Q_2-Q_1}{\dfrac{Q_1+Q_2}{2}}}{{\dfrac{P_2-P_1}{\dfrac{P_1+P_2}{2}}}}"


"\text{Price elasticity of demend}=\dfrac{\dfrac{45-55}{\dfrac{45+55}{2}}}{{\dfrac{20-12}{\dfrac{20+12}{2}}}}"


"\text{Price elasticity of demend}=\blue{-0.4}"